OpenStudy (anonymous):

2 square root of 32 minus 4 square root of 18 plus 3 square root of 128

OpenStudy (anonymous):

\[2\sqrt{32}-4\sqrt{18}+3\sqrt{128}\]

OpenStudy (anonymous):

please help me

OpenStudy (mathlegend):

I will break down one of them... and you will break down the other two based off of what I do. Which one do you want me to break down?

OpenStudy (anonymous):

i got 4 \[\sqrt{2}\]

OpenStudy (anonymous):

\[4 \sqrt{?}\]

OpenStudy (mathlegend):

How did you get that? I got something else.

OpenStudy (mathlegend):

Show me your steps please.

OpenStudy (anonymous):